# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v2

@@ Abstract
Large language models (LLMs) can exhibit bias in a variety of ways. Such biases can create or exacerbate unfair outcomes for certain groups within a protected attribute, including, but not limited to sex, race, sexual orientation, or age. This paper aims to provide a technical guide for practitioners to assess bias and fairness risks in LLM use cases. The main contribution of this work is a decision framework that allows practitioners to determine which metrics to use for a specific LLM use case. To achieve this, this study categorizes LLM bias and fairness risks, maps those risks to a taxonomy of LLM use cases, and then formally defines various metrics to assess each type of risk. As part of this work, several new bias and fairness metrics are introduced, including innovative counterfactual metrics as well as metrics based on stereotype classifiers. Instead of focusing solely on the model itself, the sensitivity of both prompt-risk and model-risk are taken into account by defining evaluations at the level of an LLM use case, characterized by a model and a population of prompts. Furthermore, because all of the evaluation metrics are calculated solely using the LLM output, the proposed framework is highly practical and easily actionable for practitioners. [^1] Note: The examples provided in this paper are purely hypothetical and are not intended to reflect the specific work or practices of the author-affiliated company. They are used solely for illustrative purposes. Any resemblance to actual practices or projects is coincidental.
@@ 1 Introduction
The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level (Anthis et al., 2024). Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023). However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024b). Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
@@
This work aims to address these limitations by developing an actionable LLM bias and fairness evaluation framework defined at the use case level. Drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values. This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024b). By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
@@
To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segments these definitions by risk category. Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible. Lastly, for each risk category, various bias and fairness evaluation metrics are detailed, with discussions provided on their input requirements, calculation methods, the risks they assess, and circumstances under which they should be applied. As part of this work, a variety of novel bias and fairness metrics are introduced. This includes innovative counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
For practical reasons, this study limits the selection of LLM bias and fairness metrics to those requiring only LLM generated output as inputs. This includes 1) generated text metrics, which take a generated set of tokens as input (Gallegos et al., 2023), 2) recommendation fairness metrics, calculated on a set of LLM-provided recommendations (Zhang et al., 2023), and 3) classification fairness metrics, which are already well-established in the machine learning fairness literature (Bellamy et al., 2018; Saleiro et al., 2018; Weerts et al., 2023; Hardt et al., 2016; Feldman et al., 2014; Mehrabi et al., 2019). Due to practical limitations of the input requirements, the framework omits both embedding-based metrics, which are computed using an LLM’s hidden vector representations of words or sentences (Islam et al., 2016; May et al., 2019; Guo and Caliskan, 2020), and probability-based metrics, which leverage predicted token probabilities from an LLM (Webster et al., 2020; Kurita et al., 2019; Ahn and Oh, 2021; Kaneko and Bollegala, 2021; Salazar et al., 2019; Barikeri et al., 2021; Nangia et al., 2020; Nadeem et al., 2021). To further support the rationale behind the selection of evaluation metrics, it is important to note that metrics focused on the downstream task, consistent with the metrics incorporated in this framework, have been shown to be more reliable than metrics derived from embeddings or token probabilities (Goldfarb-Tarrant et al., 2020; Delobelle et al., 2021).
The remainder of this paper is organized as follows. Section 2 formally defines various notions of bias and fairness in LLMs. Section 3 discusses various techniques for conducting bias and fairness assessments in LLMs. Section 4 offers a framework for choosing among LLM bias and fairness metrics based on use case characteristics and stakeholder values. Finally, Section 5 offers concluding remarks.
@@ 2 Bias and Fairness Risks for LLM Use Cases
This section formally defines several pre-requisite terms and concepts upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al. (2023). [^2] The notation and terminology used in this paper differs from those used by Gallegos et al. (2023). The approach outlined below adopts modified versions of these definitions to better fit the tone and purpose. For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023). The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others. Formal establishment of these concepts is achieved through the use of mathematical notation whenever feasible. The aim is to maintain consistency in notation throughout, even if it deviates from the notation used in the referenced papers.
@@ 2.1 Preliminary Definitions
This section provides preliminary definitions to be used throughout the subsequent sections.
@@
An LLM $\mathcal{M}:\mathcal{X}\xrightarrow{}\mathcal{Y}$ is a pre-trained, transformer-based model that maps a text sequence $X\in\mathcal{X}$ to an output $\hat{Y}\in\mathcal{Y}$, where $\mathcal{X}$ denotes the set of all possible text inputs (i.e. prompts) and the form of $\hat{Y}$ is specific to the LLM and the use case (Gallegos et al., 2023). [^3] Following Gallegos et al. (2023), $\mathcal{M}$ is assumed to have an autoregressive, autoencoding, or encoder-decoder architecture and has undergone training on an extensive corpus containing hundreds of millions to trillions of tokens.  [^4] Hereafter, all LLM inputs $X$ are referred to as ‘prompts’. Let $\mathcal{X}$ denote the set of all possible prompts. Let $\theta$ parameterize $\mathcal{M}$, such that $\hat{Y}=\mathcal{M}(X;\theta)$.
@@
A population of prompts, denoted $\mathcal{P}_{X}$, is a collection of LLM inputs. To characterize well-defined use cases, subsequent sections refer to a ‘known population of prompts’, indicating that practitioners possess information about the prompt domain and are able draw representative samples from $\mathcal{P}_{X}$. For instance, a population of prompts might consist of clinical notes, where each individual prompt includes a collection of notes, accompanied by specific instructions for the LLM to generate a summary (Chuang et al., 2024).
@@
Protected Attribute Groups.
@@
A protected attribute group lexicon $A\in\mathcal{A}$ is a collection of words that correspond to protected attribute group $G\in\mathcal{G}$ [^5] Hereafter, in a slight abuse of notation, $\mathcal{G}$ (and analogously $\mathcal{A}$) can denote either a partition of or the union of each $G\in\mathcal{G}$ (and analogously each $A\in\mathcal{A}$). [^6] One example, provided by Bommasani et al. (2023), of protected attribute words for the protected attribute group ‘males’ is { ‘he’, ‘son’, ‘his’, ‘him’, ‘father’, ‘man’, ‘boy’, ‘himself’, ‘male’, ‘brother’, ‘sons’, ‘fathers’, ‘men’, ‘boys’, ‘males’, ‘brothers’, ‘uncle’, ‘uncles’, ‘nephew’, ‘nephews’ }.  [^7] The examples that consider male vs. female are used in accordance with the convention established by previous studies such as Bordia and Bowman (2019) and Bommasani et al. (2023). It is important to note that this usage is solely for the purpose of consistency and does not intend to imply that there are only two genders. The acknowledgement of a broader spectrum of gender identities is recognized and respected.
@@ 2.2 LLM Bias and Fairness Risks
This section presents formal definitions of various notions of bias and fairness applicable to LLMs. A taxonomy for these definitions is proposed, organized by risk category. Specifically, the risk categories included in this taxonomy are toxicity, stereotyping, counterfactual fairness, and allocational harms.
@@ 2.2.1 Toxicity
Drawing from the definitions of toxicity and derogatory language outlined in the survey conducted by Gallegos et al. (2023), let the definition of toxic text encompass any offensive language that 1) launches attacks, issues threats, or incites hate or violence against a social group, or 2) includes the usage of pejorative slurs, insults, or any other forms of expression that specifically target and belittle a social group. To formalize this, a corresponding fairness desideratum is introduced below.
@@ 2.2.2 Stereotyping
Stereotyping is an important type of social bias that should be considered in the context of LLMs (Bommasani et al., 2023; Bordia and Bowman, 2019; Zekun et al., 2023). This study follows the work by Gallegos et al. (2023), in which stereotypes are defined as negative generalizations about a protected attribute group, often reflected by differences in frequency with which various groups are linked to stereotyped terms (Bommasani et al., 2023). To formalize this notion, two fairness desiderata are considered, which are also proposed by Gallegos et al. (2023): equal group associations (EGA) and equal neutral associations (ENA).
@@
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies equal group associations if, for each $w\in W,P(w\in\hat{Y}|\hat{Y}\cap A^{\prime}\neq\emptyset)=P(w\in\hat{Y}|\hat{Y}\cap A^{\prime\prime}\neq\emptyset)$. Put simply, equal group associations requires that each neutral word in $W$ is equally likely to be contained in the output of $\mathcal{M}$ regardless of which protected attribute group is mentioned.
Equal Neutral Associations (Gallegos et al., 2023).
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ with respective associated lexicons $A^{\prime},A^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies Equal Neutral Associations if, $P(A^{\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset)=P(A^{\prime\prime}\cap\hat{Y}\neq\emptyset|\hat{Y}\cap W\neq\emptyset).$ In other words, Equal Neutral Associations requires that co-occurrence of protected attribute words with a set of neutral words $W$ is equally probable for both protected attribute groups. [^9] In practice, the set $W$ referenced in the definitions of equal group associations and equal neutral associations typically contains frequently stereotyped terms such as professions or adjectives (Bordia and Bowman, 2019; Bommasani et al., 2023).
@@ 2.2.3 Counterfactual Fairness
In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024b). Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness. Depending on context and stakeholder values, the practitioner may wish to assess an LLM use case for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt. Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al. (2023).
@@
where $(X^{\prime},X^{\prime\prime})$ is a counterfactual input pair corresponding to $G^{\prime},G^{\prime\prime}$ (Gallegos et al., 2023). [^10] This modified definition of counterfactual invariance relaxes strict equality as in Gallegos et al. (2023).
@@ 2.2.4 Allocational Harms
Allocational harms, which Gallegos et al. (2023) define as an unequal distribution of resources or opportunities among different protected attribute groups, have been widely studied in the machine learning fairness literature (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Kamishima et al., 2012; Zhang et al., 2018; Hardt et al., 2016; Feldman et al., 2014; Pleiss et al., 2017; Kamiran et al., 2012; Agarwal et al., 2018; Kamiran and Calders, 2011; Chouldechova, 2016). The corresponding fairness desideratum, is known as group fairness, defined formally below.
@@ 2.3 Mapping Bias and Fairness Risks to LLM Use Cases
Section 2.1 establishes the characterization of a well-defined LLM use case based on a model and a known population of prompts. [^11] Note that use case segmentation proposed in this work is by task, which can be controlled through various means such as fine-tuning of the LLM, providing examples with few-shot prompting, or incorporating instructions in system or user prompts. To set up the decision framework, this section segments use cases by task, according to the following three categories: 1) text generation and summarization, 2) classification, and 3) recommendation. Descriptions and examples are provided in Table 1.
The bias and fairness evaluation framework proposed in this work is intended for large-scale applications, in which the volume of generated responses makes exhaustive human review impractical. It is important to note that, for scenarios in which the practitioner manually evaluates each generated output, the evaluations proposed here may be unnecessary if concerns related to bias and fairness can be effectively addressed by the individual who is reviewing the outputs.
Table 1: Taxonomy of LLM Use Cases and Associated Bias/Fairness Risks
Use Case Category Description Examples Bias/Fairness Risk
Text Generation and Summarization.

### v3

@@ Abstract
Large language models (LLMs) can exhibit bias in a variety of ways. Such biases can create or exacerbate unfair outcomes for certain groups within a protected attribute, including, but not limited to sex, race, sexual orientation, or age. In this paper, we propose a decision framework that allows practitioners to determine which bias and fairness metrics to use for a specific LLM use case. To establish the framework, we define bias and fairness risks for LLMs, map those risks to a taxonomy of LLM use cases, and then define various metrics to assess each type of risk. Instead of focusing solely on the model itself, we account for both prompt-specific- and model-specific-risk by defining evaluations at the level of an LLM use case, characterized by a model and a population of prompts. Furthermore, because all of the evaluation metrics are calculated solely using the LLM output, our proposed framework is highly practical and easily actionable for practitioners. For streamlined implementation, all evaluation metrics included in the framework are offered in this paper’s companion Python toolkit, LangFair. Finally, our experiments demonstrate substantial variation in bias and fairness across use cases, underscoring the importance of use-case-level assessments. [^1] Note: The examples provided in this paper are purely hypothetical and are not intended to reflect the specific work, practices, or opinions of the author-affiliated company. They are used solely for illustrative purposes. Any resemblance to actual practices or projects is coincidental. All opinions are the author’s own.
@@ 1 Introduction
The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) makes evaluating bias and fairness at the model level difficult (Anthis et al., 2024). Existing approaches largely rely on benchmark datasets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023). However, these assessments do not account for prompt-specific risks that have been shown to significantly influence the likelihood of biased and unfair LLM responses (Wang et al., 2024b). [^2] Wang et al. (2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts. Moreover, to the best of our knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
@@
To address these limitations, we propose an LLM bias and fairness evaluation framework defined at the use case level. Drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018), our framework enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering the task, relevant characteristics of the prompts, and stakeholder values. This evaluation approach is unique in that it follows a bring-your-own-prompts approach, in which metrics are computed from LLM responses to actual prompts from the practitioner’s use case. Our framework is designed for well-defined use cases, where prompts are sampled from a known population, allowing for bias and fairness assessments that are customized for a specific application.
@@
To introduce the framework, we first define bias and fairness risks for LLMs from the literature and map those risks to a taxonomy of use cases. For each risk category, we then present various evaluation metrics and discuss their input requirements, computation methods, the risks they assess, and circumstances under which they should be applied. As part of this work, we also introduce a variety of novel bias and fairness metrics. Specifically, these new metrics include counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well as a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics. For practical reasons, we limit the selection of LLM bias and fairness metrics to those requiring only LLM generated output for computation.
To streamline implementation of the framework, all included bias and fairness metrics are offered by this paper’s companion Python toolkit, LangFair. [^3] https://github.com/cvs-health/langfair In practice, users provide a sample of prompts from their use case and their LLM of choice, and LangFair streamlines generation of LLM responses and computes applicable metrics for their use case. This toolkit offers a model-agnostic, user-friendly way to implement our evaluation framework for real-world use cases.
Lastly, we conduct a series of experiments to evaluate bias and fairness in several text generation and summarization use cases. In particular, we construct 6 unique use cases, characterized by three sets of prompts and two LLMs. We find substantial variation in bias and fairness across use cases, underscoring the importance of use-case-level assessments.
@@ 2 Bias and Fairness Risks for LLM Use Cases
In this section, we define various terms upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al. (2023). [^4] The notation and terminology used in this paper differs from those used by Gallegos et al. (2023). We adopt modified versions of these definitions to better fit the tone and purpose. For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023a; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023). Note that, throughout this paper, we explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others.
@@ 2.1 Preliminary Definitions
Below we provide several preliminary definitions to be used throughout the subsequent sections.
@@
An LLM $\mathcal{M}:\mathcal{X}\xrightarrow{}\mathcal{Y}$ is a pre-trained, transformer-based model that maps a text sequence $X\in\mathcal{X}$ to an output $\hat{Y}\in\mathcal{Y}$, where $\mathcal{X}$ denotes the set of all possible text inputs (i.e. user prompts) and the form of $\hat{Y}$ is specific to the LLM and the use case (Gallegos et al., 2023). [^5] Following Gallegos et al. (2023), $\mathcal{M}$ is assumed to have an autoregressive, autoencoding, or encoder-decoder architecture and has undergone training on an extensive corpus containing hundreds of millions to trillions of tokens.  [^6] Hereafter, all LLM inputs $X$ are referred to as ‘prompts’. Let $\mathcal{X}$ denote the set of all possible prompts. Let $\theta$ parameterize $\mathcal{M}$, such that $\hat{Y}=\mathcal{M}(X;\theta)$.
@@
A population of prompts, denoted $\mathcal{P}_{X}$, is a collection of LLM inputs (user prompts). To characterize well-defined use cases, we subsequently refer to a ‘known population of prompts’, indicating that practitioners possess information about the prompt domain and are able draw representative samples from $\mathcal{P}_{X}$. For instance, a population of prompts might consist of clinical notes, where each individual prompt includes a collection of notes, accompanied by specific instructions for the LLM to generate a summary (Chuang et al., 2024).
@@
Protected Attribute Group.
@@
A protected attribute group lexicon $A\in\mathcal{A}$ is a collection of words that correspond to protected attribute group $G\in\mathcal{G}$ [^7] Hereafter, in a slight abuse of notation, $\mathcal{G}$ (and analogously $\mathcal{A}$) can denote either a partition of or the union of each $G\in\mathcal{G}$ (and analogously each $A\in\mathcal{A}$). [^8] One example, provided by Liang et al. (2023), of protected attribute words for the protected attribute group ‘males’ is { ‘he’, ‘son’, ‘his’, ‘him’, ‘father’, ‘man’, ‘boy’, ‘himself’, ‘male’, ‘brother’, ‘sons’, ‘fathers’, ‘men’, ‘boys’, ‘males’, ‘brothers’, ‘uncle’, ‘uncles’, ‘nephew’, ‘nephews’ }.  [^9] The examples that consider male vs. female are used in accordance with the convention established by previous studies such as Bordia and Bowman (2019) and Liang et al. (2023). It is important to note that this usage is solely for the purpose of consistency and does not intend to imply that there are only two genders. The acknowledgment of a broader spectrum of gender identities is recognized and respected.
@@ 2.2 LLM Bias and Fairness Risks
In this section, we define various bias and fairness risks applicable to LLMs and define corresponding desiderata. Namely, these risks include toxicity, stereotyping, counterfactual fairness, and allocational harms.
@@ 2.2.1 Toxicity
Following Gallegos et al. (2023), we define toxic text as any offensive language that 1) launches attacks, issues threats, or incites hate or violence against a social group, or 2) includes the usage of pejorative slurs, insults, or any other forms of expression that specifically target and belittle a social group. To formalize this, we introduce a corresponding desideratum known as non-toxicity.
@@ 2.2.2 Stereotyping
Stereotyping is an important type of social bias that should be considered in the context of LLMs (Liang et al., 2023; Bordia and Bowman, 2019; Zekun et al., 2023). We follow Gallegos et al. (2023) and define stereotypes as negative generalizations about a protected attribute group, often reflected by differences in frequency with which various groups are linked to stereotyped terms (Liang et al., 2023). The corresponding desideratum, proposed by Gallegos et al. (2023), is known as equal group associations.
@@
For two protected attribute groups $G^{\prime},G^{\prime\prime},$ and a set of neutral words $W$, an LLM use case $(\mathcal{M},\mathcal{P}_{X})$ satisfies equal group associations if, for each $w\in W,P(w\in\hat{Y}|\hat{Y}\cap A^{\prime}\neq\emptyset)=P(w\in\hat{Y}|\hat{Y}\cap A^{\prime\prime}\neq\emptyset)$. Put simply, equal group associations requires that each neutral word in $W$ is equally likely to be contained in the output of $\mathcal{M}$, regardless of which protected attribute group is mentioned.
@@ 2.2.3 Counterfactual Fairness
In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024b). Following previous work (Huang et al., 2020; Garg et al., 2019), we refer to this concept as (lack of) counterfactual fairness. Depending on context and stakeholder values, the practitioner may wish to assess an LLM use case for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt. Below, we present the corresponding fairness desideratum, known as counterfactual invariance, adapted from Gallegos et al. (2023).
@@
where $(X^{\prime},X^{\prime\prime})$ is a counterfactual input pair corresponding to $G^{\prime},G^{\prime\prime}$ (Gallegos et al., 2023). [^11] This modified definition of counterfactual invariance relaxes strict equality as in Gallegos et al. (2023). Note that some metrics that assess counterfactual invariance assess similarity, rather than dissimilarity, meaning the inequality will be reversed.
@@ 2.2.4 Allocational Harms
Allocational harms, which Gallegos et al. (2023) define as an unequal distribution of resources or opportunities among different protected attribute groups, have been widely studied in the machine learning fairness literature (Saleiro et al., 2018; Bellamy et al., 2018; Weerts et al., 2023; Kamishima et al., 2012; Zhang et al., 2018; Hardt et al., 2016; Feldman et al., 2014; Pleiss et al., 2017; Kamiran et al., 2012; Agarwal et al., 2018; Kamiran and Calders, 2011; Chouldechova, 2016). In this work, we measure allocational harms based on group fairness, defined below.
@@ 2.3 Mapping Bias and Fairness Risks to LLM Use Cases
In section 2.1, we characterize an LLM use case based on a model and a known population of prompts. Here, we categorize use cases into three task-based groups: 1) text generation and summarization, 2) classification, and 3) recommendation. [^12] Note that our case segmentation is by task, which can be controlled through various means such as fine-tuning of the LLM, providing examples with few-shot prompting, or incorporating instructions in system or user prompts. [^13] Note that our proposed framework is intended for large-scale applications, in which the volume of generated responses makes exhaustive human review impractical. For scenarios in which the practitioner manually evaluates each generated output, the evaluations we propose may be unnecessary if concerns related to bias and fairness can be effectively addressed by the individual who is reviewing the outputs. Descriptions and examples are provided in Table 1.
2.3.1 Text Generation and Summarization